The NetCDF C++ Interface Guide

Class Documentation Version 4.0.1 6 August 2008

Russ Rew Unidata Program Center

Copyright \begin{matrix}c \\ ◯ \end{matrix} 2005-2006 University Corporation for Atmospheric Research

Permission is granted to make and distribute verbatim copies of this manual provided that the copyright notice and these paragraphs are preserved on all copies. The software and any accompanying written materials are provided “as is” without warranty of any kind. UCAR expressly disclaims all warranties of any kind, either expressed or implied, including but not limited to the implied warranties of merchantability and fitness for a particular purpose. The Unidata Program Center is managed by the University Corporation for Atmospheric Research and sponsored by the National Science Foundation. Any opinions, findings, conclusions, or recommendations expressed in this publication are those of the author(s) and do not necessarily reflect the views of the National Science Foundation. Mention of any commercial company or product in this document does not constitute an endorsement by the Unidata Program Center. Unidata does not authorize any use of information from this publication for advertising or publicity purposes.

Table of Contents

Introduction . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1

Class Hierarchy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 Auxiliary Types and Constants . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

NetCDF Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Class NcFile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

Public Member Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 Class NcDim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

Public Member Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Class NcTypedComponent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6

Public Member Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 Class NcVar. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

Public Member Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 Class NcAtt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Public Member Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

Auxiliary Classes . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

Class NcValues. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

Public Member Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 Class NcError . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

Public Member Functions . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16

Index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 17

Introduction

The main requirements for the design of the C++ interface are:

• to provide the functionality of the C interface;

• to provide type safety by eliminating all use of void* pointers; and

• to provide an interface that is simpler to use than the C interface.

Some of the features of the C++ interface are:

• No IDs needed for netCDF’s variables, dimensions, or attributes.

• No explicit open or close calls needed for netCDF files; a constructor opens and a destructor closes a file.

• No need to specify types for creating attributes; they will have the type of the value provided.

• No use of void*: values are type-checked.

• Less indirection is needed for dimensions and dimension sizes than with the C interface. A variable’s dimensions can be provided as arguments when defining a variable.

• Code for data types is isolated to make the addition of new types easier.

• No explicit ncredef or ncendef calls are needed for switching between define and data modes. Whenever a mode switch is required, it happens implicitly.

The header file ‘netcdfcpp.h’ must be included in source code files using this interface.

This release provides some of the functionality of netCDF version 4, but not for the enhanced data model introduced with netCDF-4.

This manual assumes familiarity with the netCDF User’s Guide, where the concepts of netCDF dimensions, variables, and attributes are discussed.

Class Hierarchy

The class for netCDF file objects is NcFile.

The components of a netCDF file are dimensions, variables, and attributes. There is a class for each of these kinds of objects; NcDim, NcVar, and NcAtt. Variables and attributes share some common characteristics that are factored out in the abstract base

class NcTypedComponent.

An auxiliary class, NcValues, provides a type for arrays of values that are read from or written to netCDF files. Another auxiliary class, NcError, provides facilities for handling errors.

NcFile                     netCDF file
NcDim                      dimension
NcTypedComponent           abstract base class
NcVar                variable
NcAtt                attribute
NcValues                   abstract base class for array
NcValues_ncbyte      array of bytes
NcValues_char        array of characters
NcValues_short       array of shorts
NcValues_int         array of ints
NcValues_long        array of longs
NcValues_float       array of floats
NcValues_double      array of doubles
NcError                    for error handling

Auxiliary Types and Constants

The netCDF classes use several auxiliary types for arguments and return types from member

functions: NcToken, NcType, NcBool, and ncbyte.

NcToken Used for names for netCDF objects, in particular variable names, dimension names, and attribute names. Currently this is just a typedef for const char*.

NcType Used for specifying netCDF external value types. Currently this is an enumerated type with the following legitimate values: ncByte, ncChar, ncShort,

ncInt, ncLong (deprecated), ncFloat, and ncDouble.

NcBool Used for the return type of some member functions. If the member function fails, 0 is returned, otherwise some non-zero value. Currently this is just a typedef for unsigned int. It will be changed to bool when all C++ compilers support the new bool type.

ncbyte Used to declare values of type ncByte, for 8-bit integer data. (This is currently a typedef for unsigned char, but it may be changed to a typedef for signed char, so don’t depend on the underlying representation.)

NetCDF Classes

Class NcFile

NcFile is the class for netCDF files, providing methods for netCDF file operations.

Some member functions return pointers to dimensions (NcDim) or variables (NcVar). These objects are owned by the NcFile they are associated with, and will be deleted automatically by the NcFile destructor (or by the close member function, if this is called earlier than the destructor), so users should not delete these. Member functions that return pointers to attributes (NcAtt) pass ownership to the calling function; users should delete attributes when they are finished with them.

Member functions that return NcBool yield TRUE on success and FALSE on failure. Member functions that return a pointer value return a NULL pointer on failure.

This class interface hides the distinction in the C and Fortran interfaces between define mode (when dimensions, variables, or attributes are being defined or renamed), and data mode (when data values are being accessed), by automatically switching between the modes when necessary. Be aware that switching from accessing data to adding or renaming dimensions, variables and attributes can be expensive, since it may entail a copy of the data.

Public Member Functions

NcFile( const char * path, FileMode = ReadOnly, size_t *chunksizeptr = NULL,
size_t initialsize = 0, FileFormat = Classic)

The constructor creates a new netCDF file or opens an existing netCDF file. The path argument may be an OPenDAP DAP URL if DAP support is enabled. The FileMode argument can be any of ReadOnly (the default) to open an existing file for reading, Write to open an existing file for reading or writing, Replace to create a new empty file even if the named file already exists, or New to create a new file only if the named file does not already exist. The optional FileFormat argument can be any of Classic (the default),

Offset64Bits, Netcdf4, or Netcdf4Classic.

The optional chunksizeptr and initialsize tuning parameters are as described in the corresponding nc__create() function in the C interface. The constructor will not fail, but in the case of a bad path name, improper permissions, or if the file already exists and you have specified FileMode as New, no netCDF file will be created or opened. If the constructor fails to create or open a netCDF file, a subsequent call to the is_valid() member function will return False.

~NcFile( void )

Destructor. The file is closed and all resources associated with it are released, including the associated NcVar and NcDim objects. If you wish to close the file earlier, you may explicitly call the close member function; a subsequent destructor call will work properly.

NcBool close( void )

Close netCDF file earlier than it would be closed by the NcFile destructor.

NcBool is_valid( void ) const

Returns TRUE if valid netCDF file, FALSE otherwise (e.g. if constructor could not open file).

int num_dims( void ) const

Returns the number of dimensions in the netCDF file.

int num_vars( void ) const

Returns the number of variables in the netCDF file.

int num_atts( void ) const

Returns the number of global attributes in the netCDF file.

NcDim* get_dim(NcToken name) const

Get a dimension by name.

NcVar* get_var(NcToken name) const

Get a variable by name.

NcAtt* get_att(NcToken name) const

Get a global attribute by name.

NcDim* get_dim(int n) const

Get the nth dimension (beginning with the 0th).

NcVar* get_var(int n) const

Get the nth variable (beginning with the 0th).

NcAtt* get_att(int n) const

Get the nth global attribute (beginning with the 0th).

NcDim* rec_dim( void ) const

Get the unlimited dimension, if any.

The following add_member functions put the file in define mode, so could be expensive. To avoid copying of data, invoke these before writing data to variables.

NcDim* add_dim(NcToken dimname)

Add an unlimited dimension named dimname to the netCDF file.

NcDim* add_dim(NcToken dimname, long dimsize)

Add a dimension named dimname of size dimsize.

NcVar* add_var(NcToken varname, NcType type, const NcDim*, ...)

Add a variable named varname of the specified type (ncByte, ncChar, ncShort, ncInt, ncFloat, ncDouble) to the open netCDF file. The variable is defined with a shape that depends on how many dimension arguments are provided. A scalar variable would have 0 dimensions, a vector would have 1 dimension, and so on. Supply as many dimensions as needed, up to 5. If more than 5 dimensions are required, use the n-dimensional version of this member function instead.

NcVar* add_var(NcToken varname, NcType type, int ndims, const NcDim** dims)

Add a variable named varname of ndims dimensions and of the specified type. This method must be used when dealing with variables of more than 5 dimensions.

NcBool add_att(NcToken name, ncbyte val)
NcBool add_att(NcToken name, char val)
NcBool add_att(NcToken name, short val)
NcBool add_att(NcToken name, int val)
NcBool add_att(NcToken name, float val)
NcBool add_att(NcToken name, double val)

Add global scalar attributes of the specified name and with the supplied value.

NcBool add_att(NcToken name, const char* val)

Add global string-valued attribute with the specified name and C string value (terminated with a \0 character).

NcBool add_att(NcToken name, int n, const ncbyte* val)
NcBool add_att(NcToken name, int n, const char* val)
NcBool add_att(NcToken name, int n, const short* val)
NcBool add_att(NcToken name, int n, const int* val)
NcBool add_att(NcToken name, int n, const float* val)
NcBool add_att(NcToken name, int n, const double* val)

Add global vector attributes with the specified name, length, and values.

NcBool set_fill(FillMode mode = Fill)

Sets fill-mode to either NcFile::Fill or NcFile::NoFill. Default is Fill, in which case unwritten values are pre-written with appropriate type-specific or variable-specific fill values.

enum NcFile::FillMode get_fill( void ) const

Returns fill mode of the file, either NcFile::Fill or NcFile::NoFill.

enum NcFile::FileFormat get_format( void ) const

Returns format version of the file, either NcFile::Classic,

          NcFile:Offset64Bits, NcFile:Netcdf4, or NcFile::Netcdf4Classic.
NcBool sync( void )

Synchronizes file to disk. This flushes buffers so that readers of the file will see recent changes.

NcBool abort( void )

Either just closes file (if recently it has been in data mode as the result of accessing data), or backs out of the most recent sequence of changes to the file schema (dimensions, variables, and attributes).

Class NcDim

A netCDF dimension has a name and a size. Dimensions are only created and destroyed by NcFile member functions, because they cannot exist independently of an open netCDF file. Hence there are no public constructors or destructors.

Public Member Functions

NcToken name( void ) const

Returns the name of the dimension if it exists, 0 otherwise.

long size( void ) const

Returns the dimension size.

NcBool is_valid( void ) const

Returns TRUE if file and dimension are both valid, FALSE otherwise.

NcBool is_unlimited( void ) const

Returns TRUE if the dimension is the unlimited dimension, FALSE if either not a valid netCDF file, or if the dimension is not the unlimited dimension.

NcBool rename( NcToken newname )

Renames the dimension to newname.

NcBool sync( void )

If the dimension may have been renamed, make sure its name is updated.

Class NcTypedComponent

NcTypedComponent is an abstract base class for NcVar and NcAtt that captures the similarities between netCDF variables and attributes. We list here the member functions that variables and attributes inherit from NcTypedComponent, but these member functions are also documented under the NcVar and NcAtt classes for convenience.

Public Member Functions

NcToken name( void ) const

Returns the name of the variable or attribute.

NcType type( void ) const

Returns the type of the variable or attribute. The type will be one of ncByte,

ncChar, ncShort, ncInt, ncFloat, or ncDouble.
NcBool is_valid( void ) const

Returns TRUE if the component is valid, FALSE otherwise.

long num_vals( void ) const

Returns the number of values for an attribute or variable. For an attribute, this is just 1 for a scalar attribute, the number of values for a vector-valued attribute, and the number of characters for a string-valued attribute. For a variable, this is the product of the dimension sizes for all the variable’s dimensions.

NcBool rename( NcToken newname )

Renames the variable or attribute.

NcValues* values( void ) const

Returns a pointer to the block of all values for the variable or attribute. The caller is responsible for deleting this block of values when no longer needed, as well as the pointer returned by the values method. Note that this is not a good way to read selected values of a variable; use the get member function instead, to get single values or selected cross-sections of values.

ncbyte as_ncbyte( int n ) const
char as_char( int n ) const
short as_short( int n ) const
int as_int( int n ) const
nclong as_nclong( int n ) const // deprecated
long as_long( int n ) const
float as_float( int n ) const
double as_double( int n ) const
char* as_string( int n ) const

Get the n-th value of the attribute or variable. These member functions provide conversions from the value type of the variable or attribute to the specified type. If the value is out-of-range, the fill-value of the appropriate type is returned.

Class NcVar

NcVar is derived from NcTypedComponent, and represents a netCDF variable. A netCDF variable has a name, a type, a shape, zero or more attributes, and a block of values associated with it. Because variables are only associated with open netCDF files, there are no public constructors for this class. Use member functions of NcFile to get variables or add new variables.

Public Member Functions

NcToken name( void ) const

Returns the name of the variable.

NcType type( void ) const

Returns the type of the variable. The type will be one of ncByte, ncChar,

          ncShort, ncInt, ncFloat, or ncDouble.
int num_dims( void ) const

Returns number of dimensions for this variable.

NcDim* get_dim( int n ) const

Returns a pointer to the n-th dimension (starting at 0). Returns a NULLpointer if an invalid dimension is requested.

long* edges( void ) const

Returns the shape of the variable, in the form of a vector containing the sizes of the dimensions of the variable. The caller is responsible for deleting the returned edge vector when no longer needed.

int num_atts( void ) const

Returns the number of attributes attached to the variable.

NcAtt* get_att( NcToken attname ) const
NcAtt* get_att( int n ) const

The first member function returns a variable attribute by name. The second returns the n-th (starting at 0) attribute of the variable. In either case, if no such attribute has been attached to the variable, zero is returned. Attributes returned in this way belong to the caller, and hence should eventually be deleted by the caller to avoid memory leaks.

NcBool is_valid( void ) const

Returns TRUE if the variable is valid, FALSE otherwise.

long num_vals( void ) const

Returns the number of values for a variable. This is just 1 for a scalar variable, or the product of the dimension sizes for all the variable’s dimensions.

NcValues* values( void ) const

Returns a pointer to the block of all values for the variable. The caller is responsible for deleting this block of values when no longer needed. Note that this is not a good way to read selected values of a variable; use the get member function instead, to get single values or selected cross-sections of values.

NcBool put(const ncbyte* vals, long c0, long c1, long c2, long c3, long c4)
NcBool put(const char* vals, long c0, long c1, long c2, long c3, long c4)
NcBool put(const short* vals, long c0, long c1, long c2, long c3, long c4)
NcBool put(const int* vals, long c0, long c1, long c2, long c3, long c4)
NcBool put(const long* vals, long c0, long c1, long c2, long c3, long c4)
NcBool put(const float* vals, long c0, long c1, long c2, long c3, long c4)
NcBool put(const double* vals, long c0, long c1, long c2, long c3, long c4)

Write scalar or 1 to 5-dimensional arrays by providing enough arguments. Arguments are edge lengths, and their number must not exceed variable’s dimensionality. Start corner is [0,0,..., 0] by default, but may be reset using the set_cur() member function for this variable. FALSE is returned if type of values does not match type for variable. For more than 5 dimensions, use the overloaded n-dimensional form of the put member function.

NcBool put(const ncbyte* vals, const long* counts)
NcBool put(const char* vals, const long* counts)
NcBool put(const short* vals, const long* counts)
NcBool put(const int* vals, const long* counts)
NcBool put(const long* vals, const long* counts)
NcBool put(const float* vals, const long* counts)
NcBool put(const double* vals, const long* counts)

Write n-dimensional arrays, starting at [0, 0, ..., 0] by default, may be reset with set_cur(). FALSE is returned if type of values does not match type for variable.

NcBool get(ncbyte* vals, long c0, long c1, long c2, long c3, long c4) const
NcBool get(char* vals, long c0, long c1, long c2, long c3, long c4) const
NcBool get(short* vals, long c0, long c1, long c2, long c3, long c4) const
NcBool get(int* vals, long c0, long c1, long c2, long c3, long c4) const
NcBool get(long* vals, long c0, long c1, long c2, long c3, long c4) const
NcBool get(float* vals, long c0, long c1, long c2, long c3, long c4) const
NcBool get(double* vals, long c0, long c1, long c2, long c3, long c4) const

Get scalar or 1 to 5 dimensional arrays by providing enough arguments. Arguments are edge lengths, and their number must not exceed variable’s dimensionality. Start corner is [0,0,..., 0] by default, but may be reset using the set_cur() member function. FALSE is returned if type of values does not match type for variable.

NcBool get(ncbyte* vals, const long* counts) const
NcBool get(char* vals, const long* counts) const
NcBool get(short* vals, const long* counts) const
NcBool get(int* vals, const long* counts) const
NcBool get(long* vals, const long* counts) const
NcBool get(float* vals, const long* counts) const
NcBool get(double* vals, const long* counts) const

Get n-dimensional arrays, starting at [0, 0, ..., 0] by default, may be reset with set_cur() member function. FALSE is returned if type of values does not match type for variable.

NcBool set_cur(long c0=-1, long c1=-1, long c2=-1, long c3=-1, long c4=-1)
NcBool set_cur(long* cur)

Resets the starting corner to the values supplied. The first form works for a variable of dimensionality from scalar to 5 dimensions. For more than five dimensions, use the second form, in which the number of longs supplied must match the rank of the variable. The method returns FALSE if any argument is greater than the size of the corresponding dimension.

NcBool add_att( NcToken, char )
NcBool add_att( NcToken, ncbyte )
NcBool add_att( NcToken, short )
NcBool add_att( NcToken, int )
NcBool add_att( NcToken, long )
NcBool add_att( NcToken, float )
NcBool add_att( NcToken, double )
NcBool add_att( NcToken, const char* )
NcBool add_att( NcToken, int, const char* )
NcBool add_att( NcToken, int, const ncbyte* )
NcBool add_att( NcToken, int, const short* )
NcBool add_att( NcToken, int, const int* )
NcBool add_att( NcToken, int, const long* )
NcBool add_att( NcToken, int, const float* )
NcBool add_att( NcToken, int, const double* )

Add scalar or vector attribute of any type to a variable, given the name, number of values, and the vector of values. These put file in define mode, so could be expensive. To avoid the expense of copying data, add attributes to variables before writing data.

NcBool rename( NcToken newname )

Renames the variable. If variable is renamed to a longer name, this puts file in define mode, so could be expensive.

ncbyte as_ncbyte( int n ) const
char as_char( int n ) const
short as_short( int n ) const
int as_int( int n ) const
nclong as_nclong( int n ) const // deprecated
long as_long( int n ) const
float as_float( int n ) const
double as_double( int n ) const
char* as_string( int n ) const

Get the n-th value of the variable, ignoring its shape. These member functions provide conversions from the value type of the variable to the specified type. If the requested value is out-of-range, the fill-value of the appropriate type is returned.

int id( void ) const

Return the variable number. This is not needed in the C++ interface, but might be needed in calling a C-function that requires that a variable be identified by number instead of name.

NcBool sync( void )

If the variable may have been renamed, make sure its name is updated.

~NcVar( void )

Destructor.

The following member functions are intended for record variables. They will also work for non-record variables, if the first dimension is interpreted as the record dimension.

long rec_size( void )
long rec_size( NcDim* )

Return the number of values per record or the number of values per dimension slice for the specified dimension.

NcValues* get_rec( void )
NcValues* get_rec( long n )

Get the data for this variable for the current record or for the nth record.

NcValues* get_rec( NcDim* )
NcValues* get_rec( NcDim*, long n )

Get the data for this variable for the current dimension slice or for the nth dimension slice.

NcBool put_rec( const ncbyte* vals )
NcBool put_rec( const char* vals )
NcBool put_rec( const short* vals )
NcBool put_rec( const int* vals )
NcBool put_rec( const long* vals )
NcBool put_rec( const float* vals )
NcBool put_rec( const double* vals )

Put a record’s worth of data for this variable in the current record.

NcBool put_rec( NcDim*, const ncbyte* vals )
NcBool put_rec( NcDim*, const char* vals )
NcBool put_rec( NcDim*, const short* vals )
NcBool put_rec( NcDim*, const int* vals )
NcBool put_rec( NcDim*, const long* vals )
NcBool put_rec( NcDim*, const float* vals )
NcBool put_rec( NcDim*, const double* vals )

Put a dimension slice worth of data for this variable in the current dimension slice.

NcBool put_rec( const ncbyte* vals, long rec )
NcBool put_rec( const char* vals, long rec )
NcBool put_rec( const short* vals, long rec )
NcBool put_rec( const int* vals, long rec )
NcBool put_rec( const long* vals, long rec )
NcBool put_rec( const float* vals, long rec )
NcBool put_rec( const double* vals, long rec )

Put a record’s worth of data for this variable in the specified record.

NcBool put_rec( NcDim*, const ncbyte* vals, long slice )
NcBool put_rec( NcDim*, const char* vals, long slice )
NcBool put_rec( NcDim*, const short* vals, long slice )
NcBool put_rec( NcDim*, const int* vals, long slice )
NcBool put_rec( NcDim*, const long* vals, long slice )
NcBool put_rec( NcDim*, const float* vals, long slice )
NcBool put_rec( NcDim*, const double* vals, long slice )

Put a dimension slice worth of data for this variable in the specified dimension slice.

long get_index( const ncbyte* vals )
long get_index( const char* vals )
long get_index( const short* vals )
long get_index( const int* vals )
long get_index( const long* vals )
long get_index( const float* vals )
long get_index( const double* vals )

Get first record index for this variable corresponding to the specified key value(s).

long get_index( NcDim*, const ncbyte* vals )
long get_index( NcDim*, const char* vals )
long get_index( NcDim*, const short* vals )
long get_index( NcDim*, const int* vals )
long get_index( NcDim*, const long* vals )
long get_index( NcDim*, const float* vals )
long get_index( NcDim*, const double* vals )

Get first index of specified dimension for this variable corresponding to the specified key value(s).

void set_rec ( long rec )

Set the current record for this variable.

void set_rec ( NcDim*, long rec )

Set the current dimension slice for the specified dimension for this variable.

Class NcAtt

NcAtt is derived from NcTypedComponent, and represents a netCDF attribute. A netCDF attribute has a name and a type, and may be either a scalar attribute or a vector attribute. Scalar attributes have one value and vector attributes have multiple values. In addition, each attribute is attached to a specific netCDF variable or is global to an entire netCDF file. Because attributes are only associated with open netCDF files, there are no public constructors for this class. Use member functions of NcFile and NcVar to get netCDF attributes or add new attributes. Most of the useful member functions for NcAtt are inherited

from class NcTypedComponent.

Public Member Functions

NcToken name( void ) const

Returns the name of the attribute.

NcType type( void ) const

Returns the type of the attribute. The type will be one of ncByte, ncChar,

          ncShort, ncInt, ncFloat, or ncDouble.
NcBool is_valid( void ) const

Returns TRUE if the attribute is valid, FALSE otherwise.

long num_vals( void ) const

Returns the number of values for an attribute. This is just 1 for a scalar attribute, the number of values for a vector-valued attribute, and the number of characters for a string-valued attribute.

NcBool rename( NcToken newname )

Renames the attribute.

NcValues* values( void ) const

Returns a pointer to the block of all values for the attribute. The caller is responsible for deleting this block of values when no longer needed.

ncbyte as_ncbyte( int n ) const
char as_char( int n ) const
short as_short( int n ) const
int as_int( int n ) const
nclong as_nclong( int n ) const // deprecated
long as_long( int n ) const
float as_float( int n ) const
double as_double( int n ) const
char* as_string( int n ) const

Get the n-th value of the attribute. These member functions provide conversions from the value type of the attribute to the specified type. If the value is outof-range, the fill-value of the appropriate type is returned.

NcBool remove( void )

Deletes the attribute from the file and detaches it from the variable. Does not call the destructor. Subsequent calls to is_valid() will return FALSE.

~NcAtt( void )

Destructor.

Auxiliary Classes

Auxiliary classes include the abstract base class NcValues, its type-specific derived subclasses, and the error-handling class NcError.

Class NcValues

Class NcValues is an abstract base class for a block of typed values. The derived classes are

NcValues_ncbyte, NcValues_char, NcValues_short, NcValues_int, NcValues_nclong
(deprecated), and NcValues_long, NcValues_float, NcValues_double. These classes are

used as the return type of the NcTypedComponent::values() member function, for typedvalue arrays associated with variables and attributes.

Public Member Functions

NcValues( void )

Default constructor.

NcValues(NcType, long)

Constructor for a value block of the specified type and length.

~NcValues( void )

Destructor.

long num( void )

Returns the number of values in the value block.

ostream& print(ostream&) const

Used to print the comma-delimited sequence of values of the value block.

void* base( void ) const

Returns a bland pointer to the beginning of the value block.

int bytes_for_one( void ) const

Returns the number of bytes required for one value.

ncbyte as_ncbyte( int n ) const
char as_char( int n ) const
short as_short( int n ) const
int as_int( int n ) const
nclong as_nclong( int n ) const // deprecated
long as_long( int n ) const
float as_float( int n ) const
double as_double( int n ) const
char* as_string( int n ) const

Provide conversions for the nth value from the value type to a desired basic type. If the value is out of range, the default "fill-value" for the appropriate type is returned.

Class NcError

This class provides control for netCDF error handling. Declaring an NcError object temporarily changes the error-handling behavior for all netCDF classes until the NcError object is destroyed (typically by going out of scope), at which time the previous error-handling behavior is restored.

Public Member Functions

NcError( Behavior b = verbose_fatal )

The constructor saves the previous error state for restoration when the destructor is invoked, and sets a new specified state. Valid error

          states are NcError::silent_nonfatal, NcError::verbose_nonfatal,
          NcError::silent_fatal, or NcError::verbose_fatal, to control whether

error messages are output from the underlying library and whether such messages are fatal or nonfatal.

~NcError( void )

Destructor, restores previous error state.

int get_err( void )

Returns most recent error, as enumerated in ‘netcdf.h’.

Index 17

Index

~

~NcAtt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 ~NcError . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 ~NcFile. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 ~NcValues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 ~NcVar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10

A

abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 add att . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 9 add dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 add var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 as char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 as double . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 as float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 as int. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 as long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 as ncbyte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 9, 12, 15 as nclong . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 as short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 as string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7, 10, 12, 15 auxiliary types and constants . . . . . . . . . . . . . . . . . . . 2

B

base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 bytes for one . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15

C

class hierarchy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 close. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3

E

edges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7

G

get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 get att . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 7 get dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 7 get err . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 get fill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 get format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 get index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 get rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 get var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

I

id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 is unlimited . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 is valid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 6, 7, 12

N

name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 7, 12 NcAtt . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 NcAtt::~NcAtt. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 NcAtt::remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 NcBool . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 ncbyte. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 NcDim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NcDim::is unlimited . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 NcDim::is valid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 NcDim::name . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 NcDim::rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 NcDim::size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 NcDim::sync. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 NcError . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15, 16 NcError::~NcError. . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 NcError::get err . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 16 NcFile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 NcFile::~NcFile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 NcFile::abort . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NcFile::add att . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NcFile::add dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::add var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::close . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 NcFile::get att. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::get dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::get fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NcFile::get format . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NcFile::get var . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::is valid . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::NcFile . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 NcFile::num atts . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::num dims. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::num vars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::rec dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 NcFile::set fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NcFile::sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 NcToken. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 NcType . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 NcTypedComponent . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 NcTypedComponent::as char . . . . . . . . . . . . 7, 10, 12 NcTypedComponent::as double . . . . . . . . . . 7, 10, 12 NcTypedComponent::as float . . . . . . . . . . . . 7, 10, 12 NcTypedComponent::as int. . . . . . . . . . . . . . 7, 10, 12 NcTypedComponent::as long . . . . . . . . . . . . 7, 10, 12 NcTypedComponent::as ncbyte . . . . . . . . . . . 6, 9, 12 NcTypedComponent::as nclong . . . . . . . . . . 7, 10, 12 NcTypedComponent::as short . . . . . . . . . . . 7, 10, 12 NcTypedComponent::as string. . . . . . . . . . . 7, 10, 12 NcTypedComponent::is valid . . . . . . . . . . . . . 6, 7, 12 NcTypedComponent::name . . . . . . . . . . . . . . . 6, 7, 12 NcTypedComponent::num vals . . . . . . . . . . . 6, 8, 12

18 NetCDF C++ Interface Guide

NcTypedComponent::rename . . . . . . . . . . . . . 6, 9, 12 NcTypedComponent::type . . . . . . . . . . . . . . . . 6, 7, 12 NcTypedComponent::values . . . . . . . . . . . . . . 6, 8, 12 NcValues . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::~NcValues . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as char. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as double . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as float . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as long. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as ncbyte . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as nclong . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::as string . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::base . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::bytes for one. . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::NcValues. . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::num . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues::print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues char . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues double . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues float. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues int . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues long . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues ncbyte . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues nclong. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcValues short . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 NcVar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 NcVar::~NcVar . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 NcVar::add att . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 NcVar::edges . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 NcVar::get . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 NcVar::get att . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 NcVar::get dim . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 NcVar::get index . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 NcVar::get rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 NcVar::id . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 NcVar::num atts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 NcVar::num dims . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 7 NcVar::put . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 NcVar::put rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 11 NcVar::rec size . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 NcVar::set cur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 NcVar::set rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 NcVar::sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 num . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 num atts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 7 num dims . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4, 7 num vals . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 8, 12 num vars . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

P

print . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 15 put . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 8 put rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10, 11

R

rec dim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 rec size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 10 remove . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 rename . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 9, 12 requirements of C++ interface . . . . . . . . . . . . . . . . . . . 1

S

set cur . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 9 set fill . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5 set rec . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6 sync . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5, 6, 10

T

type . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 7, 12

V

values . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 6, 8, 12